# Growth of Gravitational Wave Spectrum from Sound Waves in a Universe with Generic Expansion Rate

### v1

@@ III $\Upsilon$ factor in singlet component universe
if $\tilde{q}$ is large. Then the integrand takes the form of a fraction. After omitting the linear terms associated with $H_{s}a_{s}R_{*}$, the oscillation term of its numerator $\mathcal{N}_{\rm osc}$ can be expressed as
@@
×$\times\left[P_{1}*{\rm cos}(x-z)-P_{3}*{\rm cos}(x^{\prime}-z)\right].$ (63)
@@
×$\times\left[P_{1}^{2}\frac{{\rm cos}(2x)}{2}+P_{3}^{2}\frac{{\rm cos}(2x^{\prime})}{2}-P_{1}P_{3}{\rm cos}\left(x+x^{\prime}\right)\right]\bigg{\}},$ (70)
@@
×$\times{\rm cos}\left(\frac{qR_{*}y_{-}}{c_{k}}\right)\left[\frac{P_{1}^{2}}{2}+\frac{P_{3}^{2}}{2}-P_{1}P_{3}{\rm cos}\left(x-x^{\prime}\right)\right]\bigg{\}}.$ (71)
@@
Thus, we expect its integral value just to provide the oscillatory behavior rather than the dominant term, and its integral value should be much smaller than the latter. This can be verified by using the Riemann-Lebesgue lemma if $y$ is large enough Stein and Murphy (1993). We also calculate its integral value over $\tilde{q}^{2}$ when $c_{k}=0.00079$ and $qR_{*}=10$ by numerical method and find that the corresponding value is 0.0286, -0.0035 and -0.0163 for $w=1/4$ and $y=5,10,50$, respectively, which is much smaller than the corresponding value of $\mathcal{G}_{2}$ in most intervals, see below.
The $\mathcal{G}_{2}$ could be calculated directly in the limit of $y_{-}\to 0$ by noting that the term in the brackets is $y_{+}$ independent, and can be written in the following form:
@@
×$\times\left(P_{1}^{2}+P_{3}^{2}-2P_{1}P_{3}{\rm cos}\left[\frac{3\pi(w-1)}{6w+2}\right]\right),$ (73)
@@
$\Upsilon=\frac{2\left[y^{3(w-1)/2}-1\right]}{3(w-1)}.$ (75)
Compared with the result of Hindmarsh and Hijazi (2019), the $\Upsilon$ factor is an extra factor originated from the finite duration time of sound waves. The $\Upsilon$ factor for different values of $w$ is shown in Fig.(2). Obviously, the integral value of $\mathcal{G}_{1}$ over $\tilde{q}^{2}$ is indeed much smaller than the $\Upsilon$ factor, so it is safe to omit the contribution of $\mathcal{G}_{1}$. Considering the radiation-dominated universe case ($w=\frac{1}{3}$) and the matter-dominated universe case ($w=0$), our new $\Upsilon$ factor could correctly reproduce our previous results Guo et al. (2021)
$\Upsilon_{\rm RD}$ $=1-\frac{1}{y},$ (77)
$\Upsilon_{\rm MD}$ $=\frac{2}{3}\left(1-\frac{1}{y^{3/2}}\right).$ (78)
An interesting aspect is that our derivation relies only on the conditions that the source is stationary and that the characteristic quantity $c_{k}$ is relatively small. Consequently, the resulting $\Upsilon$ factor is actually independent of the sound shell model. This suggests that $\Upsilon$ may serve as a more universal factor to describe how gravitational waves induced by a stationary source evolve over time Roper Pol et al. (2024).
We find that although $\Upsilon$ does not diverge in the vicinity of $w=1$, it behaves like a logarithmic function, suggesting that $\Upsilon$ tends to infinity when the active time of the source is sufficiently long. To understand the underlying reasons, we reexamine Eq. (III) and observe that the logarithmic behavior stems from the power of $y_{+}$ in the integrand, which equals $-1$ when $w=1$. This prompts us to analyze the composition of this power. There are three contributions to this power: (1) the Jacobian determinant provides a factor of $(y_{1}y_{2})^{\frac{1-3w}{2}}$; (2) the Green’s function contributes a factor of $y_{+}^{2}$; (3) the UETC of the energy-momentum tensor $\Pi^{2}$ introduces a factor of $y_{1}^{4}y_{2}^{4}$, which is canceled by $a^{2}(\eta_{1})a^{2}(\eta_{2})$, leaving a term $y_{1}^{2}y_{2}^{2}$, as shown in Eq.(II). Thus, we can express $\Upsilon$ roughly as
[…]
$y=\frac{a(t)}{a_{s}}=\left(\frac{t}{t_{s}}\right)^{\frac{2}{3(1+w)}}.$ (81)
Figure 3: Schematic diagram of $y$ as a function of time $t$.
IV GW spectrum today
With $\Upsilon$ obtained, the gravitational wave spectrum today can be written in the following factorized form:
$a_{s}H_{s}R_{*}\approx\left(8\pi\right)^{1/3}v_{w}\frac{H_{s}}{\beta},$ (87)
×$\times\left(\frac{a_{e}}{a_{0}}\right)^{4}\left(\frac{H_{e}}{H_{0}}\right)^{2}\frac{y^{3(w+1)}}{y^{4}}\Upsilon(y;w).$ (88)
V conclusion
In this work, we successfully utilized the sound shell model in an expanding universe to derive the factor $\Upsilon$, which reflects the growth of gravitational waves induced by sound waves in a single-component universe that satisfies the strong energy condition. We obtained the following observations: (1) Our new $\Upsilon$ accurately reproduces previous results for radiation-dominated and matter-dominated universes; (2) During the kination period, the new $\Upsilon$ exhibits a logarithmic behavior, driven by the competition between the rate of the gravitational wave production and the dilution rate of the gravitational wave source; (3) The new $\Upsilon$ does not depend on the sound shell model, suggesting that it may represent a more universal result and is not limited to cosmological phase transitions.
Acknowledgments
We would like to thank Ligong Bian and Rui-Xiao Zhang for helpful discussions. This work is supported by the National Natural Science Foundation of China (12105248, 11821505, 12075300, 12335005, 12147103), by Peng-Huan-Wu Theoretical Physics Innovation Center (12047503), and by the Research Fund for Outstanding Talents from Henan Normal University (5101029470335).

### v2

@@ Abstract
if $\tilde{q}$ is large. After simplification (see Appendix for detail), the integral with respect to $y_{+}$ can ultimately be written as
where
$\Upsilon=\frac{2\left[y^{3(w-1)/2}-1\right]}{3(w-1)}.$ (26)
Compared with the result of Hindmarsh and Hijazi (2019), the $\Upsilon$ factor is an extra factor originated from the finite duration time of sound waves. The $\Upsilon$ factor for different values of $w$ is shown in Fig.(1). Considering the radiation-dominated universe case ($w=\frac{1}{3}$) and the matter-dominated universe case ($w=0$), our new $\Upsilon$ factor could correctly reproduce our previous results Guo et al. (2021)
$\Upsilon_{\rm RD}$ $=1-\frac{1}{y},$ (28)
$\Upsilon_{\rm MD}$ $=\frac{2}{3}\left(1-\frac{1}{y^{3/2}}\right).$ (29)
An interesting aspect is that our derivation relies only on the conditions that the source is stationary and that the characteristic quantity $R_{*}a_{s}H_{s}$ is relatively small. Consequently, the resulting $\Upsilon$ factor is actually independent of the sound shell model. This suggests that $\Upsilon$ may serve as a more universal factor to describe how gravitational waves induced by a stationary source evolve over time Roper Pol et al. (2024).
We find that although $\Upsilon$ does not diverge in the vicinity of $w=1$, it behaves like a logarithmic function, suggesting that $\Upsilon$ tends to infinity when the active time of the source is sufficiently long. To understand the underlying reasons, we reexamine Eq. (.1) and observe that the logarithmic behavior stems from the power of $y_{+}$ in the integrand, which equals 1-1 when $w=1$. This prompts us to analyze the composition of this power. There are three contributions to this power: (1) the Jacobian determinant provides a factor of $(y_{1}y_{2})^{\frac{1-3w}{2}}$; (2) the Green’s function contributes a factor of $y_{+}^{2}$; (3) the UETC of the energy-momentum tensor $\Pi^{2}$ introduces a factor of $y_{1}^{4}y_{2}^{4}$, which is canceled by $a^{2}(\eta_{1})a^{2}(\eta_{2})$, leaving a term $y_{1}^{2}y_{2}^{2}$, as shown in Eq.(Growth of Gravitational Wave Spectrum from Sound Waves in a Universe with Generic Expansion Rate). Thus, we can express $\Upsilon$ roughly as
$\Upsilon$ $\sim\int dy_{+}~y_{+}^{2}\times\frac{\left(y_{1}y_{2}\right)^{\frac{1-3w}{2}}}{y_{1}^{2}y_{2}^{2}}$
$=\int dy_{+}\frac{{\rm Green^{\prime}s~function}}{{\rm Source}}\times{\rm Jacobian~Det}$
$=\int dy_{+}~\frac{y_{+}^{2}}{y_{+}^{\frac{8}{1+3w}}}y_{+}^{\frac{2-6w}{1+3w}}=\int dy_{+}~y_{+}^{\frac{-4}{1+3w}},$ (31)
$y=\frac{a(t)}{a_{s}}=\left(\frac{t}{t_{s}}\right)^{\frac{2}{3(1+w)}}.$ (32)
Figure 2: Schematic diagram of $y$ as a function of time $t$.
Spectrum Today–With $\Upsilon$ obtained, the gravitational wave spectrum today can be written in the following factorized form:
$a_{s}H_{s}R_{*}\approx\left(8\pi\right)^{1/3}v_{w}\frac{H_{s}}{\beta},$ (38)
$\times\left(\frac{a_{e}}{a_{0}}\right)^{4}\left(\frac{H_{e}}{H_{0}}\right)^{2}\frac{y^{3(w+1)}}{y^{4}}\Upsilon(y;w).$ (39)
Conclusion–In this work, we successfully utilized the sound shell model in an expanding universe to derive the factor $\Upsilon$, which reflects the growth of gravitational waves induced by sound waves in a single-component universe that satisfies the strong energy condition. We obtained the following observations: (1) Our new $\Upsilon$ accurately reproduces previous results for radiation-dominated and matter-dominated universes; (2) During the kination period, the new $\Upsilon$ exhibits a logarithmic behavior, driven by the competition between the rate of the gravitational wave production and the dilution rate of the gravitational wave source; (3) The new $\Upsilon$ does not depend on the sound shell model, suggesting that it may represent a more universal result and is not limited to cosmological phase transitions.
Acknowledgments–We would like to thank Ligong Bian and Rui-Xiao Zhang for helpful discussions. This work is supported by the National Natural Science Foundation of China (12105248, 11821505, 12075300, 12335005, 12147103), by Peng-Huan-Wu Theoretical Physics Innovation Center (12047503), and by the Research Fund for Outstanding Talents from Henan Normal University (5101029470335).
Supplemental Appendix
.1 Derivation of $\Upsilon$ factor
By substituting the asymptotic form of the Bessel function, the integral over $y_{+}$ can be simplified into a fractional expression. After omitting the linear terms associated with $H_{s}a_{s}R_{*}$, the oscillatory term of its numerator $\mathcal{N}_{\rm osc}$ can be expressed as
@@
$\times\left[P_{1}*{\rm cos}(x-z)-P_{3}*{\rm cos}(x^{\prime}-z)\right].$ (51)
@@
$\times\left[P_{1}^{2}\frac{{\rm cos}(2x)}{2}+P_{3}^{2}\frac{{\rm cos}(2x^{\prime})}{2}-P_{1}P_{3}{\rm cos}\left(x+x^{\prime}\right)\right]\bigg{\}},$ (58)
@@
$\times{\rm cos}\left(\frac{qR_{*}y_{-}}{c_{k}}\right)\left[\frac{P_{1}^{2}}{2}+\frac{P_{3}^{2}}{2}-P_{1}P_{3}{\rm cos}\left(x-x^{\prime}\right)\right]\bigg{\}}.$ (59)
@@
Thus, we expect its integral value just to provide the oscillatory behavior rather than the dominant term, and its integral value should be much smaller than the latter. This can be verified by using the Riemann-Lebesgue lemma if $y$ is large enough Stein and Murphy (1993). We also calculate its integral value over $\tilde{q}^{2}$ when $c_{k}=0.00079$ and $qR_{*}=10$ by numerical method and find that the corresponding value is 0.0286, -0.0035 and -0.0163 for $w=1/4$ and $y=5,10,50$, respectively, which is much smaller than the corresponding value of $\mathcal{G}_{2}$ in most intervals. The $\mathcal{G}_{2}$ could be calculated directly in the limit of $y_{-}\to 0$ by noting that the term in the brackets is $y_{+}$ independent, and can be written in the following form:
@@
$\times\left(P_{1}^{2}+P_{3}^{2}-2P_{1}P_{3}{\rm cos}\left[\frac{3\pi(w-1)}{6w+2}\right]\right),$ (61)